OpenStudy (anonymous):

What is the slope for the following table of values? gallons of gas used 1 2 3 4 miles driven 26 52 78 104 2 4 30 26

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) slope=(52-26)/(2-1) slope=26